OpenStudy (anonymous):

how do I find the y and x intercepts? "y= -2x +1"

OpenStudy (jhannybean):

just use the replace and solve method. Replace y with 0, solve for x. : x intercept. Replace x with 0, solve for y : y-intercept

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

y= -2x +1 0= -2x + 1 -1 -1 -1 = -2x

OpenStudy (anonymous):

The problem is I don't know what to do once I am at this part...

OpenStudy (jhannybean):

in order to isolate x, we have to divide both sides by -2

jimthompson5910 (jim_thompson5910):

agreed with Jhannybean -2x means "-2 times x" so if we had -2x = -1, and we wanted to isolate x, we undo the multiplication of the -2 the opposite of multiplication is division. You must divide both sides by -2 to isolate x

OpenStudy (anonymous):

You first have to divide -2x by both sides so you would end up with y= x + -2

OpenStudy (anonymous):

so x + 0.5?

jimthompson5910 (jim_thompson5910):

I think you meant `x = 0.5` ?

OpenStudy (anonymous):

oops I mean x = 0.5

jimthompson5910 (jim_thompson5910):

correct. That or x = 1/2 which is the same thing in different form

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

@Jhannybean why do that? You stated above that y = 0. That's what led to x = 0.5 Don't you mean plug in x = 0 ?

OpenStudy (anonymous):

so I plot one on (0.5,0) and another on (0,1) ?

OpenStudy (jhannybean):

That's true...i thought you could do either or. inputting it into my calculator i realize that it isn't possible.

OpenStudy (anonymous):

was what I put correct?

jimthompson5910 (jim_thompson5910):

`so I plot one on (0.5,0) and another on (0,1) ?` correct @Amineh0923 now draw a straight line through those points

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910 thanks

jimthompson5910 (jim_thompson5910):

no problem